# Convexified Message-Passing Graph Neural Networks

### v1

@@ 6 Conclusions and Future Work
We presented Convexified Message-Passing Graph Neural Networks (CGNNs), a novel framework that transforms the training of message-passing GNNs to a convex optimization problem. They ensure global optimality and computational efficiency, while having provable generalization and statistical guarantees. Empirically, we demonstrated that our CGNNs framework can effectively capture complex patterns without the need for over-parameterization or extensive data. This insight offers a more efficient and data-scarce alternative to traditional deep learning approaches, with substantial improvements across various benchmark datasets. As discussed in Section 4.4, the training of CGNNs is heavily affected by the factorization of the kernel matrix in terms of time and space complexity. Addressing these challenges is thus a key direction for future work.
Appendix A Upper Bounding the Nuclear Norm
@@
Appendix B Choice of Kernels and Valid Activations
In this section, we provide properties regarding two types of kernels: the inverse polynomial kernel (Section B.1) and the Gaussian RBF kernel (Section B.2). We show that the reproducing kernel Hilbert space (RKHS) corresponding to these kernels contain the nonlinear filter at each layer $\ell$ for the $f^{\text{th}}$ feature of node $i$ for certain choices of an activation function $\sigma$. Finally, we summarize the possible choices for an activation function $\sigma$ that result from our discussion (Section B.3).
B.1 Inverse Polynomial Kernel
@@
Zhang et al. [59] prove that $\kappa^{IP}$ indeed constitutes a legal kernel function. We include the proof to make the paper self-contained: they provide a feature mapping $\varphi:\mathbb{R}^{F}\rightarrow\ell^{2}(\mathbb{N})$ for which $\kappa^{IP}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi(\mathbf{z}),\varphi(\mathbf{z}^{\prime})\rangle$. The $(k_{1},\dots,k_{j})$-th coordinate of $\varphi(\mathbf{z})$, where $j\in\mathbb{N}$ and $k_{1},\dots,k_{j}\in[F_{\ell}]$, is given by $2^{-\frac{j+1}{2}}x_{k_{1}}\dots x_{k_{j}}$. Thus:
@@
as desired. Combining (7) with the proof of Lemma 1 in Appendix A.1 of Zhang et al. [58], the RKHS associated with $\kappa^{IP}$ is comprised of the class of nonlinear graph filters in (7). It can be formulated as follows:
Corollary B.1.
@@
Consequently, Zhang et al. [58] state that upper bounding $C_{\sigma}(t)$ for a particular activation function $\sigma$ is sufficient. To enforce $C_{\sigma}(t)<\infty$, the scalars $\{a_{j}\}_{j=0}^{\infty}$ must rapidly converge to zero, implying that $\sigma$ must be sufficiently smooth. For polynomial functions of degree $d$, the definition of $C_{\sigma}$ yields $C_{\sigma}(t)=\mathcal{O}(t^{d})$. For the sinusoid activation $\sigma(t):=sin(t)$, we have:
@@
[…]
Remark B.2.
@@
Regarding Corollary B.1, as well as Corollary B.3 which follows in the next subsection, for enforcing $C_{\sigma}(t)<\infty$, the coefficients $\{a_{j}\}_{j=0}^{\infty}$ in $\sigma$’s polynomial expansion must quickly converge to zero as illustrated above, thus requiring $\sigma$ to be sufficiently smooth. Such activations include polynomial functions, the sinusoid, the erf function and the smoothed hinge loss.
B.2 Gaussian RBF Kernel
@@
Combining (7) with the proof of Lemma 2 in Appendix A.2 of Zhang et al. [58], the RKHS associated with $\kappa^{RBF}$ is comprised of the class of nonlinear input features. It can be formulated as follows:
Corollary B.3.
@@
Comparing Corollary B.1 and Corollary B.3, $\kappa^{RBF}$ requires a stronger condition on the smoothness of the activation function. For polynomial functions of degree $d$, $C_{\sigma}(t)=\mathcal{O}(t^{d})$ remains intact. However, for the sinusoid activation $\sigma(t):=\sin(t)$, we have:
@@
[…]
B.3 Valid Activation Functions

### v2

@@ 6 CONCLUSION
We presented CGNNs, which transform the training of message-passing GNNs to a convex optimization problem. They ensure global optimality and computational efficiency, while having provable generalization and statistical guarantees. Empirically, we showed that our CGNNs can effectively capture complex patterns without the need for over-parameterization or extensive data. This insight offers a more efficient and data-scarce alternative to traditional deep learning approaches, with substantial improvements across various benchmark datasets. As noted in Section 4.4, CGNNs’ training is heavily affected by the factorization of the kernel matrix in terms of time and space complexity. Addressing these challenges is thus a key direction for future work.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes. Our mathematical setting and model are described in Section 3, while in Sections 4.24.5 we clarify that certain theoretical claims, such as inclusion in the RKHS and generalization guarantees in Theorem 4.7, hold only for certain choices of kernel functions and sufficiently smooth activation functions (see Appendices C.1-C.3). Our algorithm for learning multi-layer CGCNs is fully described in Section 4.3.]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes. Section 4.4 and Appendix F discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No. The paper fully discloses all the information needed to reproduce the main experimental results. Further, all code is publicly available at (Cohen et al., 2026) for reproducibility as well as further inspection.]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes. In Sections 4.24.5, we clarify that certain theoretical claims, such as inclusion in the RKHS and generalization guarantees in Theorem 4.7, hold only for certain choices of kernel functions (e.g., Gaussian RBF) and sufficiently smooth activation functions.]
    2. (b)
      Complete proofs of all theoretical results. [Yes. Lemma 4.1 is proven in Appendix B, in Appendix D we prove Lemma 4.2, Lemma 4.3 in proven in Appendix E, and Theorem 4.7 is discussed in Section 4.5 along with a proof sketch, while a complete proof appears in Appendix G.]
    3. (c)
      Clear explanations of any assumptions. [Yes. See Appendices C.1-C.2 for possible choices of kernels and Appendix C.3 for valid activation functions, as well as an elaboration on sufficient smoothness in Remark C.2.]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [No. The paper fully discloses all the information needed to reproduce the main experimental results. All code is publicly available at (Cohen et al., 2026) for reproducibility as well as further inspection.]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes. In Section 5 and Appendix H.2, we specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.). The statistics of the considered datasets appear in Appendix H.1, while Appendix H.2 includes additional implementation details.]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes. In Table 1 we explicitly report standard deviations. Similarly, the bar plots in Figure 3 report mean accuracies over four runs along with error bars that denote the standard deviation.]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes. The paper provides sufficient information on the computing resources needed to reproduce the experiments in Appendix H.2.]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes. The creators and original owners of assets used in the paper are properly credited.]
    2. (b)
      The license information of the assets, if applicable. [Yes. The license and terms of use are explicitly mentioned and properly respected. We also state which version of the asset is used and, where possible, include a URL.]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable. The paper does not release new assets.]
    4. (d)
      Information about consent from data providers/curators. [Yes. As mentioned earlier, the license and terms of use are explicitly mentioned and properly respected.]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable.]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable. The paper does not involve crowdsourcing nor research with human subjects.]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable. The paper does not involve crowdsourcing nor research with human subjects.]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable. The paper does not involve crowdsourcing nor research with human subjects.]
Convexified Message-Passing Graph Neural Networks:
Supplementary Materials
Appendix A ADDITIONAL RELATED WORK
A.1 Graph Transformers
Instead of aggregating local neighborhoods, Graph Transformers capture interaction information between any pair of nodes via a single self-attention layer. Some works design specific attention mechanisms or positional encodings (Hussain et al., 2022; Kreuzer et al., 2021; Ma et al., 2023; Ying et al., 2021), while others combine message-passing GNNs to build hybrid architectures (Chen et al., 2022; Rampášek et al., 2022; Shi et al., 2021; Wu et al., 2021). Those methods enable nodes to interact with any other node, facilitating direct modeling of long-range relations. Our framework broadly applies to such hybrid methods, as we exhibit on the leading models GraphGPS (Rampášek et al., 2022) and GraphTrans (Shi et al., 2021; Wu et al., 2021). Empirically, applying our framework consistently improves performance in graph classification tasks, with significant gains on GraphGPS, showcasing the versatility and effectiveness of our approach.
A.2 Comparison to Convexified CNNs (Zhang et al., 2017)
There is a fundamental mathematical gap between CNNs (Zhang et al., 2017) and message-passing GNNs studied in our work. The transition from grids to graphs is not merely an application of existing theory; it requires solving a distinct set of theoretical problems related, which Zhang et al. (2017) do not address. Our work thus differs from that by Zhang et al. (2017) in several fundamental dimensions, such as:
  1. 1.
    Irregular Domains: Zhang et al. (2017) rely on the fixed, grid-like structure of images where ”neighbors” are deterministic; the input has a constant dimension and regular structure. They cannot account for a domain where the ”receptive field” (neighborhood) size varies from node to node as in GNNs. Further, Zhang et al. (2017) cannot handle information exchange between nodes. In contrast, our work focuses on irregular domains consisting of graphs with varying numbers of nodes and edges, which cannot be derived from the fixed-grid analysis of Zhang et al. This lack of stationarity breaks many simplifications used in the CNN setting.
  2. 2.
    Differences in Convexification Procedures: Zhang et al. (2017) use RKHS constructions tied to convolutional filters and patch extraction on grids. The kernel acts on image patches and the analysis exploits fixed patch geometry. Conversely, our work requires a different RKHS construction designed to explicitly model the recursive aggregation of information across multiple hops.
  3. 3.
    Nuclear norm relaxation: Unlike Zhang et al. (2017), our nuclear norm relaxation is applied to matrix representations of graph filters that concatenate hop-wise filters across features. The low-rank structure thus has node-hop-feature semantics (not just patches in a grid).
Finally, we show how the convexification can be applied to a wide family of message-passing GNNs and hybrid transformer models, which cannot be treated by Zhang et al. (2017).
A.3 Fundamental Distinctions from Cohen and Agmon (2021a)
While our work and that by Cohen and Agmon (2021a) use convexification in graph contexts, the architectural foundations, scope and mathematical formulations are fundamentally different.
  • Modeling and Scope: Cohen and Agmon (2021a) proposed convex versions of specific GNN models termed as aggregation GNN, which apply only to a signal with temporal structure; particularly, their formulation is restricted to a fixed communication topology. Their work is also narrowly tailored and focused on a specific application in swarm robotics, where convexity arises from modeling assumptions tailored to domain adaptation. In contrast, our CGNN framework establishes a general and theoretically principled convexification of diverse message-passing GNNs by transforming the standard architecture itself, independent of application-specific constraints. In other words, we resolve key mathematical limitations of their work:
    1. 1.
      Cohen and Agmon (2021a): Their formulation is mathematically restricted to a fixed communication topology. Consequently, their method is undefined if the number of nodes changes between training and inference. It cannot be applied to standard graph classification benchmarks.
    2. 2.
      Our CGNN Framework: In contrast, our framework is explicitly designed for multiple, variable-sized graphs. By relaxing the message-passing filters directly (rather than polynomial coefficients of a fixed operator), our method supports training on multiple, heterogeneous graphs. As detailed in Appendix H, our implementation handles datasets with varying numbers of nodes and edges, a capability that is mathematically impossible under the formulation of Cohen and Agmon (2021a).
  • Modularity and Generalization: Unlike Cohen and Agmon (2021a), whose formulation does not preserve standard GNN modularity and is limited to a single fixed graph and a specific control setting, our framework retains the usual GNN architectural structure and supports multiple graphs during training. Our convexification also extends to, e.g., node-level tasks by applying the same procedure with a node-level loss on the final-layer features; no further reformulation is required. In contrast, their model cannot be directly adapted to generic node or graph prediction without substantial redesign. Hence, CGNNs provide a scalable, domain-agnostic convexification framework with a fundamentally different mathematical basis. In summary, the key modularity differences are:
    1. 1.
      Cohen and Agmon (2021a): Narrowly tailored to a specific control setting (swarm robotics) and requires substantial redesign to adapt to other tasks.
    2. 2.
      Our Work: Our framework preserves the modularity of standard GNNs; extending our convexification to node-level tasks (as opposed to graph-level) requires only a change in the final loss function, not a reformulation of the architecture.
  • Tighter Theoretical Foundations and Different Models: Our nuclear norm relaxation step and derived upper bound on relaxed filters’ nuclear norm are fundamentally different, much tighter and more general than those presented by Cohen and Agmon (2021a).
    1. 1.
      The analysis by Cohen and Agmon (2021a) is limited to polynomial filters. After the nuclear norm relaxation of polynomial filters, Cohen and Agmon (2021a) prove that the nuclear norm of the filters at layer $\ell$ is at most $R_{\ell}F_{\ell}F_{\ell-1}\sqrt{\sum_{i}(\sum_{k}\lambda_{i}^{k})^{2}}$. Here, $\lambda_{1},\dots,\lambda_{n}$ are the distinct eigenvalues of the shift operator $\mathcal{S}$, $R_{\ell},F_{\ell},K$ are as in Section 3.
    2. 2.
      Our analysis is fundamentally different: it is not restricted to polynomial filters, but applies to general filters. After the nuclear norm relaxation, Lemma 4.1 gives an upper bound on the filters’ nuclear norm, whose derivation is simpler and yields a much tighter upper bound for general filters. Namely, we prove that the nuclear norm of any possible filter at layer $\ell$ is at most $R_{\ell}\sqrt{F_{\ell}F_{\ell-1}(K+1)}$. Eventually, this results in a smaller generalization gap bound, as observed in Theorem 4.7.
  • Algorithmic Differences: Hence, our relaxation to a reproducing kernel Hilbert space (RKHS) and its analysis also substantially differ from Cohen and Agmon (2021a).
    1. 1.
      The methods of Cohen and Agmon (2021a) are limited to learning a relaxation of the frequency representation of polynomial filters. Further, their algorithm is restricted to using classic projected gradient descent instead of more modern optimizers. Also, the implementation of their framework is confined to graphs with the same number of nodes.
    2. 2.
      As before, our analysis is not restricted to polynomial filters. Rather, it applies to general non-convex filters, by directly learning the relaxed filters themselves. Our algorithm can also employ a projected version of any modern optimizer (e.g., Adam). Additionally, as detailed in Appendix H, our implementation utilizes several components that allow use to train our convex models on datasets with graphs containing varying numbers of nodes and edges.
    In short, the algorithmic differences are:
    1. 1.
      Cohen and Agmon (2021a): Restricted to learning a relaxation of the frequency representation of polynomial filters using classic projected gradient descent.
    2. 2.
      Our Work: Our relaxation to a Reproducing Kernel Hilbert Space (RKHS) allows us to directly learn the relaxed filters using any modern optimizer (e.g., Adam), ensuring better scalability and convergence in practice.
A.4 GNNs as Unified Optimization Problems
Our convexification framework relates to and fundamentally differs from prior work on GNNs as unified optimization problems. Chen et al. (2023b) first classify spatial GNNs based on how they aggregate neighbor information, while discussing how graph convolution can be depicted as an optimization problem. They also discuss how filters of spectral GNNs can be approximated via either linear, polynomial or rational approximation. The works by He et al. (2021), Wang and Zhang (2022) and Guo and Wei (2023) focus on polynomial approximation, treating spectral filters as parameterized by a polynomial basis. Specifically:
  • BernNet (He et al., 2021) parameterizes an arbitrary spectral filter in a Bernstein polynomial basis. BernNet is analyzed from the perspective of graph optimization, showing that any polynomial filter that attempts to approximate a valid filter has to take the form of BernNet.
  • Wang and Zhang (2022) notice that spectral GNNs with different polynomial bases of the spectral filters have the same expressive power but different empirical performance. Thus, for **linear** GNNs, they analyze the effect of polynomial basis from an optimization perspective. Specifically, they show that a set of orthonormal bases whose weight function corresponds to the graph signal density can enable linear GNNs to achieve the highest convergence rate.
  • Guo and Wei (2023) continue that work, by first proposing for learning an orthonormal polynomial basis. Then, they tackle the definition of optimal basis from an optimization perspective in [3], proposing a model that computes the optimal basis for a given graph structure and graph signal.
However, in all those works training remains non-convex. Our contribution is thus orthogonal and complementary: rather than deriving architectures, we convexify the training problem itself for general, non-convex message-passing GNNs (not only spectral ones). Training thus becomes a convex optimization problem, ensuring global optimality as well as an accurate and rigorous analysis of statistical properties. Further, our framework applies directly to general message-passing GNNs and hybrid transformer architectures like GraphGPS and GraphTrans, whereas The works by He et al. (2021), Wang and Zhang (2022) and Guo and Wei (2023) only cover spectral GNNs. In fact, the optimization-oriented analysis of Wang and Zhang (2022) is restricted to linear GNNs, while we focus on convexifying non-convex GNNs. We will add a dedicated discussion explicitly contrasting these settings.
A.5 Classic Graph Kernels
  1. 1.
    The relation to Weisfeiler-Lehman: The most direct link lies in the Weisfeiler-Lehman (WL) graph kernel (Shervashidze et al., 2011). It is well-established that standard message-passing GNNs are functionally equivalent to the 1-WL isomorphism test in terms of expressive power (Xu et al., 2019). Classical WL kernels operate by explicitly mapping graphs to high-dimensional sparse feature vectors containing counts of color-refined subtrees, which are then typically optimized via Support Vector Machines (SVMs).
    • The WL kernel approach: The kernel method separates feature extraction (WL refinement) from learning (SVM). The ”learning” is convex (quadratic programming), but the features are fixed and discrete.
    • Our Approach: Our CGNN framework internalizes this process. By mapping the nonlinear message-passing filters themselves into an RKHS, we do not merely extract fixed features; we learn the propagation mechanism itself. Crucially, we achieve this while retaining the hallmark benefit of the SVM era: convexity. CGNNs thus guarantee global optima, eliminating the local minima issues plaguing standard non-convex GNN training.
  2. 2.
    Implicit vs. Explicit Infinite-Dimensional Spaces: Classical graph kernels rely on the kernel trick to implicitly operate in high-dimensional spaces without computing coordinates. Conversely, widely-used GNNs (e.g., GCN, GIN) operate in explicit, low-dimensional Euclidean spaces. Convexified GNNs occupy a middle ground that leverages the best of both worlds. We utilize the RKHS formalism to first parameterize graph filters in a countable-dimensional feature space. By the representer theorem, we give a reduction of the original ERM problem to a finite-dimensional one. Then, we solve the learning problem via projected gradient descent. This mirrors the Infinite-Width GNN (or Neural Tangent Kernel) regime, but with a tractable, finite-data implementation. While Graph Neural Tangent Kernels (GNTK) (Du et al., 2019; Krishnagopal and Ruiz, 2023) are often static and used as fixed priors, CGNNs remain adaptable, learning the representation within the convex constraint set.
  3. 3.
    Dataset Implications and Generalization: The connection is empirically relevant given the standard evaluation benchmarks. Datasets such as TUDataset (e.g., MUTAG, PROTEINS, NCI1) were originally curated to benchmark graph kernels. For years, kernel methods (like the WL-subtree kernel and Deep Graph Kernels) outperformed early GNNs on these tasks due to the small sample sizes, where over-parameterized GNNs were prone to overfitting. By convexifying GNNs, our method introduces rigorous regularization inherent to the RKHS norm. This explains the strong performance of CGNNs on these “kernel-dominated” datasets: we provide the adaptivity of a neural network but with the complexity control and generalization guarantees typically associated with SVM-based kernel methods. Thus, CGNNs can be viewed as the natural neural evolution of the WL-kernel: preserving the global optimization guarantees while enabling end-to-end learning of continuous message-passing dynamics.
Appendix B UPPER BOUNDING THE NUCLEAR NORM
@@
Appendix C CHOICE OF KERNELS AND VALID ACTIVATIONS
In this section, we provide properties regarding two types of kernels: the inverse polynomial kernel (Section C.1) and the Gaussian RBF kernel (Section C.2). We show that the reproducing kernel Hilbert space (RKHS) corresponding to these kernels contain the nonlinear filter at each layer $\ell$ for the $f^{\text{th}}$ feature of node $i$ for certain choices of an activation function $\sigma$. Finally, we summarize the possible choices for an activation function $\sigma$ that result from our discussion (Section C.3).
C.1 Inverse Polynomial Kernel
@@
Zhang et al. (2017) prove that $\kappa^{IP}$ indeed constitutes a legal kernel function. We include the proof to make the paper self-contained: they provide a feature mapping $\varphi:\mathbb{R}^{F}\rightarrow\ell^{2}(\mathbb{N})$ for which $\kappa^{IP}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi(\mathbf{z}),\varphi(\mathbf{z}^{\prime})\rangle$. The $(k_{1},\dots,k_{j})$-th coordinate of $\varphi(\mathbf{z})$, where $j\in\mathbb{N}$ and $k_{1},\dots,k_{j}\in[F_{\ell}]$, is given by $2^{-\frac{j+1}{2}}x_{k_{1}}\dots x_{k_{j}}$. Thus:
@@
as desired. Combining (7) with the proof of Lemma 1 in Appendix A.1 of Zhang et al. (Zhang et al., 2016a), the RKHS associated with $\kappa^{IP}$ is comprised of the class of nonlinear graph filters in (7). It can be formulated as follows:
Corollary C.1.
@@
Consequently, Zhang et al. (Zhang et al., 2016a) state that upper bounding $C_{\sigma}(t)$ for a particular activation function $\sigma$ is sufficient. To enforce $C_{\sigma}(t)<\infty$, the scalars $\{a_{j}\}_{j=0}^{\infty}$ must rapidly converge to zero, implying that $\sigma$ must be sufficiently smooth. For polynomial functions of degree $d$, the definition of $C_{\sigma}$ yields $C_{\sigma}(t)=\mathcal{O}(t^{d})$. For the sinusoid activation $\sigma(t):=sin(t)$, we have:
@@
For the erf function $\sigma_{erf}(t):=\frac{2}{\sqrt{\pi}}\int_{0}^{t}e^{z^{2}}dz$ and the smoothed hinge loss function $\sigma_{sh}(t):=\int_{-\infty}^{t}\frac{1}{2}(\sigma_{erf}(t)+1)dz$, Zhang et al. (Zhang et al., 2016a) proved that $C_{\sigma}(t)=\mathcal{O}(e^{ct^{2}})$ for universal numerical constant $c>0$.
Remark C.2.
@@
Regarding Corollary C.1, as well as Corollary C.3 which follows in the next subsection, for enforcing $C_{\sigma}(t)<\infty$, the coefficients $\{a_{j}\}_{j=0}^{\infty}$ in $\sigma$’s polynomial expansion must quickly converge to zero as illustrated above, thus requiring $\sigma$ to be sufficiently smooth. Such activations include polynomial functions, the sinusoid, the erf function and the smoothed hinge loss.
C.2 Gaussian RBF Kernel
@@
Combining (7) with the proof of Lemma 2 in Appendix A.2 of Zhang et al. (Zhang et al., 2016a), the RKHS associated with $\kappa^{RBF}$ is comprised of the class of nonlinear input features. It can be formulated as follows:
Corollary C.3.
@@
Comparing Corollary C.1 and Corollary C.3, $\kappa^{RBF}$ requires a stronger condition on the smoothness of the activation function. For polynomial functions of degree $d$, $C_{\sigma}(t)=\mathcal{O}(t^{d})$ remains intact. However, for the sinusoid activation $\sigma(t):=\sin(t)$, we have:
@@
In contrast, for both the erf function $\sigma_{erf}(t):=\frac{2}{\sqrt{\pi}}\int_{0}^{t}e^{z^{2}}dz$ and the smoothed hinge loss function $\sigma_{sh}(t):=\int_{-\infty}^{t}\frac{1}{2}(\sigma_{erf}(t)+1)dz$, $C_{\sigma}(t)$ is infinite. Hence, the RKHS induced by $\kappa^{RBF}$ does not contain input features activated by those two functions.
C.3 Valid Activation Functions